OpenStudy (anonymous):

What is the equation of a line passing through (3,4) and having a slope of -5?

OpenStudy (muzzack):

use the point slope form \[\huge\ y-4=-5(x-3)\]

OpenStudy (anonymous):

y-4=-5(x-3)

OpenStudy (muzzack):

\[\huge\cal\ y-4=-5x-15\]

OpenStudy (anonymous):

y=-5x-11

OpenStudy (muzzack):

\(\huge\cal\ welcome~to~openstudy\)

OpenStudy (muzzack):

\[\huge\cal\ y=-5x-11\] sorry

OpenStudy (anonymous):

Answer choices: y=5x+19 y=-5x+11 y=-5x+19 y=-1/5x+19 y=-1/5x+11

OpenStudy (anonymous):

Sorry we were wrong the correct answer is the second choice. \[y=-5x+11\] @HOOP_G0d

OpenStudy (muzzack):

oh wait now i know

OpenStudy (muzzack):

when we multipplie -5*-3 it becomes a positive number

OpenStudy (muzzack):

so the answer is A

OpenStudy (muzzack):

i mean C

OpenStudy (anonymous):

You mean B?